OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution?

OpenStudy (anonymous):

0.4(12)=4.8L of alcohol now This means 7.2 L of water. (7.2+x)=0.7(12+x) 7.2+x=8.4+0.7x 0.3x=1.2 x=4 x is the amount of water added.

OpenStudy (anonymous):

Fool can i get your help

OpenStudy (anonymous):

Sure tinkerbell, how can I help you?

OpenStudy (anonymous):

im trying to check 20 math questions and see if i have the right answers

OpenStudy (anonymous):

20 is a big number :(

OpenStudy (anonymous):

i know but it will go quick i promise

OpenStudy (lalaly):

lol fool has time:P

OpenStudy (anonymous):

i have it done im just checking to see if i have them right

OpenStudy (anonymous):

lol, lala and diya would certainly help you :P

OpenStudy (anonymous):

okay sure tinker post them.

OpenStudy (lalaly):

hehehe

OpenStudy (anonymous):

post them here or do a new post

OpenStudy (anonymous):

ok